Find an equation for the line satisfying the given conditions
Through (2,8); perpendicular to 9x+7y=74
possible answers
7x+9y=-58 ans.
9x-7y=-58
7x-9y=-58
7x-9y=1
Through (-3,-2); perpendicular to -7x+9y=-42
possible answers
9x+7y=-41 ans.
9x-7y=-42
-7x-9y=-41
9x-7y=-41

Answer by checkley77(12844)

Through (2,8); perpendicular to 9x+7y=74
7y=-9x+74
y=-9x/7+74/7
A perpendicular line would have a slope=7/9
8=7/9*2+b
8=14/9+b
b=8-14/9
b=(8*9-14)/9
b=(72-14)/9
b=58/9
y=7x/9+58/9
9y=7x+58
7x-9y=-58 ans.
-------------------------------------------------
Through (-3,-2); perpendicular to -7x+9y=-42
9y=7x-42
y=7x/9-42/9 (red line)
A perpendicular line has a slope=-9/7
-2=-9/7*-3+b
-2=27/7+b
b=-2-27/7
b=(-2*7-27)/7
b=(-14-27)/7
b=-41/7
y=-9x/7-41/7
7y=-9x-41
9x+7y=-41 ans.